OpenStudy (anonymous):

Solve the proportion x/x^2-2 = 2/3x +5 What is the value of x?

OpenStudy (ghazi):

go for the cross multiplication...

OpenStudy (ghazi):

x(3x+5)=2(x^2-2).....3x^2+5x=2x^2-4...now rearrange and solve for x

OpenStudy (anonymous):

How would you rearrange that?

OpenStudy (ghazi):

3x^2-2x^2+5x+4=0

OpenStudy (ghazi):

that means X^2+5X+4=0 therefore (x+4)(x+1)=0...now you can find x

OpenStudy (anonymous):

Okay, thanks :)